solve completely the differential equation d^2y/dx^2-a^2y=0
1
Expert's answer
2019-04-25T10:59:34-0400

Solution

y''-a2y=0

Looking for a solution in the form y=eλx

(eλx)''-a2eλx=0

λ2eλx-a2eλx=0

λ2-a2=0

λ=a or λ=-a

if a not equal to zero

y1=c1eax and y2=c2e-ax

y(x)=c1eax+c2e-ax

if a=0

y1=c1 and y2=c2x

y(x)=c1+c2x

Answer

y(x)=c1eax+c2e-ax , if a not equal to zero

y(x)=c1+c2x , if a=0